OpenStudy (thats_kyy):

Paul plans to put concrete on a rectangular portion of his driveway. The portion is 12 feet long and 6 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Paul needs is $108.89. Which of the following is closest to the width of the portion of the driveway on which Paul plans to put concrete? [1 foot = 12 inches; 1 yard = 3 feet] 3 feet 5 feet 8 feet 10 feet

OpenStudy (thats_kyy):

@loch

OpenStudy (thats_kyy):

@lochana

OpenStudy (anonymous):

Actually its B: 5 feet

OpenStudy (lochana):

yes. answer is B. but you need to explain. @Summeriscool

OpenStudy (anonymous):

1.11=(4)×(x)×(16) 1.11=23x x=1.11×32 x≈1.67 1.67×3≈5.01

OpenStudy (anonymous):

sleek-feathered one

OpenStudy (anonymous):

n i gga

OpenStudy (lochana):



OpenStudy (lochana):

volume of concrete = x.0.5.12 feet^3

OpenStudy (lochana):

price of cubic concrete for 1 cubic feet = $98/3^2

OpenStudy (lochana):

\[\frac{98}{27}.x.\frac{1}{2}.12 = 108.89\]

OpenStudy (anonymous):

yo chill bruh